OpenStudy (anonymous):

Distribution with fractions. -2(6 - 9/2 x)

OpenStudy (anonymous):

Do you know how to remove brackets in this type of questions?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Nw let's take a simple example. 2(3-2) Let's first remove the bracket to solve this. 2(3-2)=2*3-2*2. Removing the brackets means to multiply the terms inside the bracket by the one outside. See if that makes a little sense

OpenStudy (anonymous):

Could you do that with the one up their ^ ? :)

OpenStudy (anonymous):

-2(6-(9/2)x)=-2*6-2*(-9/2)x Nw do u get t?

OpenStudy (anonymous):

@Lepgirl

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

No. I didnt simplify t yet.

OpenStudy (anonymous):

I hav removed the bracket. see f u cn simplify t. If u cnt tel me.

OpenStudy (anonymous):

I can't figure out the answer after simplifying . ): I'm not good at this .

OpenStudy (anonymous):

what do u get when u multiply -2 by 6?

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

what do u get when u multiply -9x/2 by -2

OpenStudy (anonymous):

i don't know how to do the 9x/2 part

OpenStudy (anonymous):

wait 9x/2 would that be 18x?

OpenStudy (anonymous):